40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 2835     Entry time: Thu Dec 16 17:07:24 2021
Author: aaron 
Type: DailyProgress 
Category: Laser 
Subject: beat note setup 

confirming fiber path

I traced the fiber beam path and confirmed that the system is set up as in elog 2833. I add a FC/PC-to-FC/APC patch cable between the 50-50 beamsplitter and the 1611FC-AC photodiode. We've been meaning to add that patch for a while to avoid damaging the connection.

I'd like to also see a beat note so we can hunt for this 140 kHz oscillation tomorrow.

I opened the connector housings on the cable for the N laser TEC and readjusted them such that the internal clamp bites down on the jacket of the cable, rather than closing over the conductor wires themselves (see attachments 1 and 2). Wiggling the cable no longer causes an alarm on the TEC.

I noticed that the FC 1611 photodiode was being powered from a GlobTek +-12 V DC supply with 0.25A current limit; however, those photodiodes require +-15 V DC with 0.3A current limit. I located the proprietary cable and am now powering that photodiode from the New Focus 0901 power supply located on the PSOMA rack.

I also noticed that the SMA cable from the fiber box panel to the spectrum analyzer that should have been carrying the beat note was connected to the wrong port! That's probably why it wasn't visible before.

With those changes, I could see the beat note on the spectrum analyzer yes.

I wanted to add a cable from the beat note PD DC output to the fiber box panel, so I could check that the DC power on that photodiode makes sense. But, I couldn't locate one in cryo or EE. Maybe the 40m or another WB lab has a spare?

Attachment 1: loose.jpg  1.891 MB  Uploaded Fri Dec 17 14:22:45 2021  | Hide | Hide all
loose.jpg
Attachment 2: clamped.jpg  1.910 MB  Uploaded Fri Dec 17 14:22:51 2021  | Hide | Hide all
clamped.jpg
ELOG V3.1.3-